OpenStudy (anonymous):

My sister has math problems. Super easy! She's in 7th grade. Can anyone help her?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

What are the problems??

OpenStudy (bossvideogamer21):

i can

OpenStudy (anonymous):

Convert the fraction to a decimal. 5/10 A. 0.56 B. 0.83 C. 5.6 D. 8.3

OpenStudy (anonymous):

Well 5/10 is .5 so the closest thing would have to be A

OpenStudy (anonymous):

Convert the fraction to a decimal. A. 11.01875 B. 11.0316 C. 11.1875 D. 11.316

OpenStudy (anonymous):

the fraction is 11 3/16

OpenStudy (anonymous):

3/16 = .1875 + 11 = 11.1875

OpenStudy (anonymous):

Which mixed number is equivalent to this decimal? 8.00027

OpenStudy (anonymous):

@fainteckos

OpenStudy (anonymous):

are there choices or will anything work?

OpenStudy (anonymous):

8 27/100,000... 8 27/10,000......8 27/1,000.....8 27/100

OpenStudy (bossvideogamer21):

800027/100000

OpenStudy (anonymous):

Thats wrong.

OpenStudy (bossvideogamer21):

and 8 27/100000

OpenStudy (anonymous):

well you cant just remove zeros or it will change its corresponding decimal

OpenStudy (bossvideogamer21):

and no its not wrong i searched it on decimal to fraction calculator

OpenStudy (anonymous):

sorry i meant 8 27/100000

OpenStudy (anonymous):

Which shows the numbers in order from least to greatest? 7 4/5.... 7.6.....7.41..

OpenStudy (anonymous):

7.41, 7.6, 7 4/5

OpenStudy (bossvideogamer21):

@EmoMuch?Yes. why does he get a medal i answered that question first

OpenStudy (anonymous):

Because he answered my first question.

OpenStudy (anonymous):

Which mixed number is greater than 8.45? A. 8 3/8 B. 8 4/9 C. 8 3/5 D. 8 5/12

OpenStudy (anonymous):

8 3/5

OpenStudy (anonymous):

A. 18 B. 36 C. 72 D. 216

OpenStudy (anonymous):

What is the least common denominator for 5/12 and 7/18

OpenStudy (anonymous):

36.

OpenStudy (anonymous):

what fraction is equivalent to 5 3/4? a. 19/4 b. 23/4 c. 15/4 d. 8/4

OpenStudy (anonymous):

B

OpenStudy (kweenlexis):

she in the 7th grade?

OpenStudy (anonymous):

Yes.

OpenStudy (kweenlexis):

im in the 7th too but im learning something different!....Does she have any more questions?

OpenStudy (anonymous):

Yes. But it takes a minute because i have to type out all of them.

OpenStudy (kweenlexis):

ok

OpenStudy (anonymous):

Which mixed number is equivalent to 904/100? a. 9 2/5 b. 9 4/10 c. 9 4/25 d. 9 1/25

OpenStudy (kweenlexis):

The answer is B.

OpenStudy (anonymous):

This next one i cant type or paste so let me figure it out 1 sec.

OpenStudy (kweenlexis):

ok

OpenStudy (anonymous):

Ugh. I give up. she said she will just guess the next one.

OpenStudy (anonymous):

Sally made the conjecture that the sum of any two prime numbers is a prime number. Which equation is a counterexample to her conjecture? A. 2 + 5 = 7 B. 15 + 4 = 19 C. 2 + 11 = 13 D. 3 + 11 = 14

OpenStudy (kweenlexis):

ok

OpenStudy (anonymous):

Anyone know the answer to that one?

OpenStudy (kweenlexis):

i do im going to explain it for her to though.

OpenStudy (anonymous):

Okay

OpenStudy (kweenlexis):

Prime numbers are numbers that are divisible by the number itself and the number 1. Prime numbers go on forever, but here is a list of prime numbers up to 20: 2, 3, 5, 7,11,13,17,19 So we are looking for an answer that makes the conjecture not true. Looking at the first answer, we see that when we add two prime numbers we get a number that is not on our list of prime numbers. So this answer shows that the conjecture is not true. This answer is the correct answer. Looking at the other answers, we see that when we add two prime numbers we get a prime number. So this answer doesn't show the conjecture is not true. So this answer is no good. The third answer again shows that when we add two prime numbers we get another prime number. So this answer doesn't show the conjecture is not true. So this answer is no good. The final answer adds a prime number and a composite number. But the conjecture says we must add two prime numbers. So this answer is no good. So the first answer is correct. 

OpenStudy (kweenlexis):

sorry its so long :(

OpenStudy (anonymous):

She said woah and her eyes were huge lol

OpenStudy (kweenlexis):

lol :) i dont blame her lol i would have done the same thing :)

OpenStudy (anonymous):

She got more than half wrong.

OpenStudy (anonymous):

Now she has to retake it.

OpenStudy (kweenlexis):

:O how did that happen??

OpenStudy (anonymous):

The last one was D.

OpenStudy (kweenlexis):

ooooopppppppsssssss SORRRRRRRYYYYYYYY :((((((((

OpenStudy (anonymous):

She said its okay. Can you help her with a few on her new test?

OpenStudy (kweenlexis):

yeah i can try lol i will tell yall if i dont know it and i may add someone i think know it :) SO BRING IT ON

OpenStudy (anonymous):

OKay.

OpenStudy (anonymous):

What is the LCD for the fractions 3/16 and 11/24?

OpenStudy (kweenlexis):

give me a second :)

OpenStudy (kweenlexis):

LCD right

OpenStudy (kweenlexis):

Rewriting input as fractions if necessary: 3/16, 11/24 For (16, 24) the least common multiple (LCM) is 48. Therefore, the least common denominator (LCD) is 48. Rewriting as equivalent fractions with the LCD: 9/48, 22/48

OpenStudy (anonymous):

48?

OpenStudy (kweenlexis):

oh sheesh i was doing the wrong thing uhhh let me redo it ok dont put that

OpenStudy (anonymous):

Okay.

OpenStudy (kweenlexis):

yeah that answer would be 48 because 16 x 3=48 and 24 x 2=48 so it would be 48. :)

OpenStudy (anonymous):

What fraction is equivalent to 4 2/3? a. 13/12 b. 26/12 c. 13/6 d. 28/6

OpenStudy (kweenlexis):

give me a secondran out of paper :)

OpenStudy (kweenlexis):

got to go get some more from down stairs

OpenStudy (anonymous):

Okay.